Quadratic equations are one of my weakest links in math which is my worst subject. Can someone assist me with these problems, I definately appreciate it!
1. Determine whether the following equations have a solution or not? Justify your answer.
s2 - 4s + 4 = 0
5/6x2 - 7x - 6/5 = 0
7a2 + 8a + 2 = 0

Answer by oscargut(2103)

if f(x)=ax^2+bx+c then
f has a solution if d=b^2-4ac >=0
s^2 - 4s + 4 = 0
d= (-4)^2-4(1)(4)=0 (has a solution with multiplicity 2)
5/6x2 - 7x - 6/5 = 0
d= (-7)^2-4(5/6)(-6/5)=53 (has 2 solutions)
7a2 + 8a + 2 = 0
d= 8^2-4(7)(2)=8 (has 2 solutions)